# Are Image Distributions Indistinguishable to Humans Indistinguishable to Classifiers?

### v1

@@ Abstract
The ultimate goal of generative models is to characterize the data distribution perfectly. For image generation, common metrics of visual quality (e.g., FID), and the truthlikeness of generated images to the human eyes seem to suggest that we are close to achieving it. However, through distribution classification tasks, we find that, in the eyes of classifiers parameterized by neural networks, the strongest diffusion models are still far from this goal. Specifically, classifiers consistently and effortlessly distinguish between real and generated images in various settings. Further, we observe an intriguing discrepancy: classifiers can identify differences between diffusion models with similar performance (e.g., U-ViT-H vs. DiT-XL), but struggle to differentiate between the smallest and largest models in the same family (e.g., EDM2-XS vs. EDM2-XXL), whereas humans exhibit the opposite tendency. As an explanation, our comprehensive empirical study suggests that, unlike humans, classifiers tend to classify images through edge and high-frequency components. We believe that our methodology can serve as a probe to understand how generative models work and inspire further thought on how existing models can be improved and how the abuse of such models can be prevented.
@@ 1 Introduction
Diffusion probabilistic models [1, 2, 3] have emerged as leading generative models for image [4, 5, 6, 7], video [8, 9, 10, 11] and 3D content generation [12, 13, 14, 15], surpassing previous generative models such as generative adversarial networks [16], variational autoencoders [17], and normalizing flows [18]. In particular, focusing on image generation, diffusion models [19, 20, 21, 22, 23, 24] can generate realistic images that are indistinguishable from real ones to the human eyes (see Fig. 2). Moreover, common quantitative metrics like FID [25] indicate that the generated distribution from state-of-the-art diffusion models is already close to the ImageNet validation set [19, 21, 20, 26, 27].
In this context, we are curious about a fundamental question: How far is the generated distribution from the real distribution? Binary classification is a natural approach, as classifiers parameterized by neural networks can serve as effective indicators of the differences between the real and generated distributions through classification accuracy. Inspired by the methodology in the work [28, 29] on dataset bias [30], we propose distribution classification tasks, predicting the distribution from which an image is taken, naturally serving as a probe to explore this critical question.
We first check are image distributions indistinguishable to humans indistinguishable to classifiers? In particular, we train commonly used classifiers to distinguish between real and generated images across various settings. To our surprise, all classifiers effortlessly achieve extremely high distribution classification accuracy with far fewer training samples than required for the diffusion models. Quantitatively, most classifiers can achieve over 98% accuracy. Furthermore, even self-supervised classifiers demonstrate the ability to identify generated images. Therefore, in the eyes of classifiers parameterized by neural networks, the generated distribution remains significantly different from the real distribution, indicating that there is still substantial room for improvement in diffusion models.
More surprisingly, we observe intriguing contradictions between classifiers and both FID and human judgments through the following two groups of experiments: The first group aims to distinguish between different diffusion models with similar FID scores and visual quality but different inductive biases. The second group aims to classify the smallest and largest models within the same diffusion model family, which share inductive biases but differ in visual quality. Notably, classifiers perform well on the first task but struggle with the second, while humans exhibit a different trend. This demonstrates that, unlike humans, classifiers focus more on inductive biases than visual quality.
@@
Furthermore, we delve into how classifiers distinguish between different distributions to better understand their behaviours. Using Grad-CAM [31], we observe that classifiers primarily rely on edge area to distinguish real and generated images. When we center-crop images at various resolutions to remove edges, classifiers still successfully differentiate based on the new edges, even at the smallest resolution of 16. Additionally, random cropping at resolutions of 128 and 64, each performed four times, maintains high classification accuracy, indicating pervasive biases recognized by classifiers. Additionally, we observe that classifiers achieve high accuracy except when only minimal low-frequency components are present. In contrast, for models within the same diffusion family, given any specific frequency bands, classifiers’ accuracy drops close to random guessing.
Our findings have several implications. Firstly, our methodology can serve as a probe to understand how generative models work. For instance, our results indicate that current diffusion models have significant room for improvement in edge areas and high-frequency components. Enhancing these aspects will better match the data distribution. Secondly, our findings suggest that the visual quality of existing diffusion models is already sufficient for content generation, as high-frequency components are generally imperceptible to humans. Lastly, our work highlights the potential for designing more efficient algorithms based on edge areas or high-frequency components to detect generated images.
@@
  • We find that classifiers can effortlessly distinguish diffusion-generated and real distributions, even if diffusion models achieve remarkably low FID and generate lifelike images (see Fig. 2).
  • We demonstrate that classifiers, unlike humans, are more sensitive to the inductive biases of different diffusion models but less sensitive to visual quality.
  • We explore how classifiers distinguish between real and generated images, finding that they primarily rely on edge information. They maintain high accuracy even with limited frequency components, except when only minimal low-frequency components are available.
@@ 2.1 Diffusion model
Diffusion models [1, 2, 3] gradually inject noise into data $\bm{x}$ during the forward process:
@@ 2.2 Frequency filter
The frequency content of an image reflects the rate of pixel value changes. Low-frequency components represent overall outline and slow grayscale changes, while high-frequency components capture fine details like object edges and textures. Notably, convolutional neural networks can capture high-frequency components that are usually imperceptible to humans [32].
@@ 3 Related work
Dataset bias. Torralba and Efros [30] introduced the "Name That Dataset" game to discover dataset bias. Tommasi et al. [29] and Liu and He [28] expanded on this by applying convolutional neural networks and large-scale datasets. Inspired by the methodologies in these works, we propose distribution classification tasks, predicting the distribution from which an image is taken. In this paper, the distribution refers to either the real distribution or the diffusion-generated distribution.
Generated image detection. We mention some empirical work on generated image detection in three key aspects: universal detection [33, 34, 35, 36], reconstruction-based methods [37, 38], and frequency analysis [32, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Frequency analysis, closely related to our work, shows that generative data usually differ from real data in the frequency spectrum, particularly in high-frequency components. Although diffusion models exhibit fewer frequency artifacts compared to GANs, they still mismatch real images in high-frequency components. This finding aligns with our results. Unlike these works, which rely on frequency spectrum analysis, our approach uses distribution classification tasks to explore the differences.
@@ 4 Experimental settings
We present the main experiment settings as follows. For more details, please see Appendix A.
Dataset. For label-to-image tasks, we use the CIFAR-10 [48] and ImageNet [49] datasets, and for text-to-image tasks, we use the COCO2014 dataset [50]. For CIFAR-10, we construct the real distribution from its training and validation sets, generating 50k training images and 10k validation images using a diffusion model for the generated distribution. For ImageNet, we consider resolutions of 256 and 512. We randomly sample 100k training images and 50k validation images from ImageNet’s respective sets for the real distribution, and generate an equivalent number of images using diffusion models for the generated distribution. We follow ADM’s [51] method to preprocess ImageNet into these resolutions. For COCO2014, we randomly sample 10k training images and 1k validation images from its respective sets for the real distribution, and use the corresponding prompts to generate an equivalent number of images for the generated distribution.
Classifier. By default, we use ResNet-50 [52] as the classifier architecture, and also consider ConvNeXt-T [53] and ViT-S [54] for completeness. Our pre-processing follows standard supervised training [53]. Specifically, during training, the classifiers process randomly augmented crops of $224\times 224$ images. For validation, images are resized to 256 pixels on the shorter side, preserving the aspect ratio, and then center cropped to $224\times 224$. For CIFAR-10 experiments, we initialize the classifier with ResNet-50 pre-trained on ImageNet and also provide the results trained from scratch (see Appendix C.1). For ImageNet, we train ResNet-50 from scratch.
Diffusion model. For the generation of CIFAR-10, we consider two diffusion models: EDM [26] and U-ViT [20]. For the generation of ImageNet-256, we explore three diffusion models: EDM2 [19], U-ViT-H/2 [20], DiT-XL/2 [21]. We resize EDM2’s generated images from 512 to 256 resolution, resulting in EDM2-XXL achieving an FID of 2.14, similar to U-ViT-H/2 and DiT-XL/2. For the generation of ImageNet-512, we use EDM2-XXL [19]. For the generation of COCO, we consider three text-to-image diffusion models: Pixart-$\alpha$ [23], SDXL [24], and Playground-v2.5 [22].
@@
5 Generated images are easily classified as fake
State-of-the-art diffusion models achieve remarkably low FID scores and can generate lifelike images (see Fig. 2). These results suggest that diffusion models are approaching the ability to perfectly learn the distribution of real images. Given this context, our initial series of experiments aims to determine whether classifiers can distinguish diffusion-generated and real distribution.
@@ 5.1 Classifiers achieve high accuracy across various settings
Label-to-image. As shown in Tab. 1, we present binary distribution classification accuracy on various combinations of real and generated distributions in the label-to-image scenario. To our surprise, in the eyes of classifiers defined by the neural network, the generated distribution is still significantly different from the real distribution. Specifically, all classifiers are capable of effortlessly distinguishing between real and generated images across various dataset combinations. Quantitatively, most classifiers can achieve over $>98\%$ distribution classification accuracy.
Moreover, we find a positive correlation between distribution classification accuracy and the number of training samples per distribution of the combination of U-ViT-H/2 and ImageNet-256 (see Fig. 2). Notably, using only 50k samples per distribution (less than 10% of the number of ImageNet samples, which the generator is trained on), classifiers achieve an accuracy exceeding 99.5%. Besides, the clear positive correlation suggests that classifiers are learning more generalizable features as the sample size increases rather than simply memorizing samples.
@@
Table 2: Four-way distribution classification on text-to-image. All classifiers yield high accuracy in distinguishing four distributions: COCO [50], Pixart-$\alpha$ [23], SDXL [24], and Playground-v2.5 [22], using only 5k or 10k training samples per dataset. Notably, as the number of training samples increases, the accuracy of distribution classification consistently improves.
@@
Figure 2: Positive correlation between accuracy and number of training samples per distribution. With only 50k samples per distribution, classifiers achieve over 99.5% accuracy.
Text-to-image. As shown in Tab. 2, we present the distribution classification accuracy in the text-to-image scenario. We consider combinations of four distributions: COCO [50], Pixart-$\alpha$ [23], SDXL [24], and Playground-v2.5 [22]. In this scenario, classifiers are tasked not only with identifying real versus generated distribution but also with determining which specific generated distribution an image comes from. Notably, we find that in all settings, classifiers achieve $>76\%$ distribution classification accuracy. Meanwhile, we also observe significant improvements as the number of training samples increases from 5k to 10k, which aligns with the label-to-image scenario.
5.2 Self-supervised classifiers can identify the generated images
Except for the supervised learning protocol, we also explore the self-supervised learning protocol. Specifically, we use the self-supervised pre-trained models, such as MAE [55] and MoCo v3 [56], to extract image features and then train a linear classifier on these features. Following, we use the self-supervised classifier to refer to this linear classifier. We report the distribution classification accuracy of self-supervised classifiers in Tab. 3. We observe two intriguing phenomena: First, distinct from semantic classification results on ImageNet, in the distribution classification task, masked image modelling-based MAE is more effective at identifying generated images related to ImageNet compared to the contrastive learning-based MoCo v3. Second, the self-supervised classifiers achieve relatively high accuracy in tasks related to ImageNet. However, there is a notable decline in accuracy when these models are transferred to CIFAR-10. We hypothesize that this occurs for two main reasons: Firstly, MAE and MoCo v3 are pre-trained specifically on ImageNet, not on CIFAR-10. Secondly, diffusion models exhibit high performance on CIFAR-10 image generation, and the mere use of a linear classifier is insufficient for distinguishing between real and generated images.
Table 3: Linear classifier on self-supervised features achieves high accuracy on ImageNet-related combinations, where the model is pre-trained, but accuracy declines when transferred to CIFAR-10. We use ViT-B as the backbone.

### v2

# Are Images Indistinguishable to Humans Also Indistinguishable to Classifiers?

@@ Abstract
The ultimate goal of generative models is to perfectly capture the data distribution. For image generation, common metrics of visual quality (e.g., FID) and the perceived truthfulness of generated images seem to suggest that we are nearing this goal. However, through distribution classification tasks, we reveal that, from the perspective of neural network-based classifiers, even advanced diffusion models are still far from this goal. Specifically, classifiers are able to consistently and effortlessly distinguish real images from generated ones across various settings. Moreover, we uncover an intriguing discrepancy: classifiers can easily differentiate between diffusion models with comparable performance (e.g., U-ViT-H vs. DiT-XL), but struggle to distinguish between models within the same family but of different scales (e.g., EDM2-XS vs. EDM2-XXL). Our methodology carries several important implications. First, it naturally serves as a diagnostic tool for diffusion models by analyzing specific features of generated data. Second, it sheds light on the model autophagy disorder and offers insights into the use of generated data: augmenting real data with generated data is more effective than replacing it.
@@ 1 Introduction
Diffusion probabilistic models (Sohl-Dickstein et al., 2015; Song et al., 2020; Ho et al., 2020) have emerged as leading generative models for image (Ramesh et al., 2021; Esser et al., 2024), video (Brooks et al., 2024; Bao et al., 2024; Zhao et al., 2024) and 3D content generation (Poole et al., 2022), surpassing previous generative models such as generative adversarial networks (Goodfellow et al., 2014), variational autoencoders (Kingma & Welling, 2013), and normalizing flows (Rezende & Mohamed, 2015). In particular, focusing on image generation, diffusion models are capable of producing realistic images that are often indistinguishable from real ones to the human eye (see Fig. 2). Furthermore, common metrics such as FID (Heusel et al., 2017) suggest that the distribution generated by state-of-the-art diffusion models closely approximates the ImageNet validation set (Karras et al., 2023; 2022; Tian et al., 2024). Therefore some works, such as Wang et al. (2024); You et al. (2024), have utilized generated data to augment training datasets. However, to the best of our knowledge, few studies have been conducted to analyze the differences between the generated and real distributions.
This raises a fundamental question: How far is the generated distribution from the real distribution? Inspired by the methodology in the work (Liu & He, 2024; Tommasi et al., 2017) on dataset bias (Torralba & Efros, 2011), we propose distribution classification tasks (see Sec. 3 for details), predicting the distribution an image belongs to - real or generated - thereby serving as a probe to explore this critical question. Notably, binary classification is a natural approach, as classifiers parameterized by neural networks can serve as effective indicators of the differences between the real and generated distributions through classification accuracy (see more analysis in Appendix  D). Compared to the widely used FID metric, our classifier-based approach directly measures distribution differences without relying on Gaussian assumptions (see further discussion in Appendix E).
We first explore are images indistinguishable to humans also indistinguishable to classifiers? To this end, we train several commonly used classifiers to distinguish between the real and generated distribution across various settings. Quantitatively, most classifiers can achieve over 98% accuracy. Meanwhile, all classifiers reach extremely high accuracy with significantly fewer training samples than the requirement of training diffusion models. Interestingly, even self-supervised classifiers exhibit this capacity to identify the generated distribution. These results demonstrate that from the perspective of neural network-based classifiers, the generated distribution remains markedly different from the real one, indicating that diffusion models still have significant room for improvement.
@@
Furthermore, we observe intriguing contradictions between classifiers’ prediction and both FID and human judgments through two groups of experiments. The first group involves classifying the smallest and largest models within the same diffusion family, which share inductive biases but differ in visual quality. The second group examines diffusion models with similar FID scores and visual quality but different inductive biases. Surprisingly, classifiers perform well on the second task but struggle with the first, revealing contradictions between classifier performance and FID scores. Meanwhile, human judgments align with classifiers in the first task but degrade to random guessing in the second, indicating inconsistency between classifier and human judgments.
Our approach displays several important implications. First, it naturally serves as a diagnostic tool for diffusion models by extending the analysis to specific features of generated data. For example, in evaluating the advanced diffusion transformer U-ViT-H/2 (Bao et al., 2023), we focus on two key aspects: spatial features and frequency ranges. Spatially, classifiers can easily distinguish whether an image is real or generated using any part of the image. In terms of frequency, classifiers struggle with minimal low-frequency information but perform well across all other frequency intervals. This suggests that U-ViT-H/2 has effectively learned low-frequency features, making generated images harder to distinguish in these regions, but remains less proficient in higher frequency regions. Second, our approach enables the evaluation of using generated data for training models. It provides a reasonable explanation for the model autophagy disorder (MAD) phenomenon observed in Shumailov et al. (2024); Alemohammad et al. (2024). This collapse is surprising because, despite the realistic appearance of the generated data, using it in a continuous loop to train subsequent generations of models leads to failure. We explain this by showing that there exists a distribution mismatch between the training and generated distributions in each generation. As this mismatch accumulates, the generated data distribution drifts further from the original. In downstream tasks, such as supervised and semi-supervised classification, this mismatch leads to performance degradation when real data is simply replaced by generated data. However, this distribution drift can be mitigated by augmenting real data with generated data, which introduces novel features while maintaining alignment with the original data distribution, leading to improved performance.
@@
  • We show that classifiers easily distinguish between diffusion-generated and real distributions, despite diffusion models achieving low FID scores and generating lifelike images (see Fig. 2).
  • We reveal the intriguing contradictions between classifier performance and widely used evaluation metrics, such as FID and human judgments.
  • We demonstrate that our approach complements traditional metrics like FID and can serve as a diagnostic tool to provide deeper insights into diffusion models.
  • We provide a reasonable explanation for the surprising model autophagy disorder phenomenon and show that augmenting real data with generated data is more effective than replacing it in supervised and semi-supervised learning.
@@ 2.1 Diffusion model
Diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020; Ho et al., 2020) gradually inject noise into data $\bm{x}$ during the forward process:
@@ 2.2 Frequency filter
The frequency content of an image represents the rate of pixel value changes. Low-frequency components capture overall shapes and gradual grayscale changes, while high-frequency components reflect fine details like edges and textures. Notably, convolutional neural networks can detect high-frequencies that are often imperceptible to humans (Wang et al., 2020a). In our paper, we implement a rectangular filter and discuss the choice between using rectangular or circular filters in Appendix F.
@@ 3 Related work
Dataset bias. Torralba & Efros (2011) introduced the "Name That Dataset" game to discover dataset bias. Tommasi et al. (2017) and Liu & He (2024) expanded on this by applying convolutional neural networks and large-scale datasets, respectively. Existing work on dataset bias focuses on classification to highlight biases in data collection for building unbiased datasets, our work leverages distribution classification to explore a fundamental question: How far is the generated distribution from the real distribution? Besides, we focus on the discrepancy between the generated distribution and the real distribution rather than between specific datasets. Furthermore, the conclusion drawn by related work highlights the ability of neural networks to detect biases within increasingly general datasets, our conclusion underscores the significant gap that still exists between current diffusion models and the real data distribution. We also highlight the potential of distribution classification methods for gaining deeper insights into generative models and more effectively leveraging generated data.
Generated image detection. Several empirical works on generated image detection span three key areas: universal detection (Liu et al., 2024; Ojha et al., 2023; Wang et al., 2020b; Girish et al., 2021), reconstruction-based methods (Wang et al., 2023b; Luo et al., 2024), and frequency analysis (Wang et al., 2020a; Zhang et al., 2019; Durall et al., 2020; Chandrasegaran et al., 2021; Frank et al., 2020; Dzanic et al., 2020; Yang et al., 2023; Corvi et al., 2023b; Ricker et al., 2022; Corvi et al., 2023a). Existing work generally focuses on either building universal discriminators to detect generated images or identifying frequency artifacts to build classifiers. In contrast, our approach leverages distribution classification to gain deeper insights into generative models and quantify the discrepancy between generated and real distributions. Instead of prioritizing classifier performance, we focus on the insights provided into generative model behavior.
@@ 4 Experimental settings
We present the main experimental settings as follows. For more details, please see Appendix A.
Dataset. For label-to-image tasks, we use the CIFAR-10 (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009) datasets, and for text-to-image tasks, we utilize the COCO2014 dataset (Lin et al., 2014). For CIFAR-10, the real distribution is constructed from its training and testing sets, comprising 50k training images and 10k testing images. The generated distribution is created using diffusion models to produce an equivalent number of images. For ImageNet, we consider resolutions of 256 and 512 and follow ADM’s (Dhariwal & Nichol, 2021) method to preprocess ImageNet into the desired resolutions. We randomly sample 100k training images and 50k validation images from the corresponding ImageNet sets to represent the real distribution, and generate an equivalent number of images using diffusion models for the generated distribution. For COCO2014, we randomly sample 10k training images and 1k validation images from the respective sets to form the real distribution, and use the corresponding prompts to generate an equivalent number of data for the generated distribution.
Classifier. By default, we use ResNet-50 (He et al., 2016) as the classifier architecture, and also consider ConvNeXt-T (Liu et al., 2022) and ViT-S (Dosovitskiy et al., 2020) for completeness. Our pre-processing follows standard supervised training (Liu et al., 2022). Specifically, during training, the classifiers process randomly augmented crops of $224\times 224$ images. For validation, images are resized to 256 pixels on the shorter side, preserving the aspect ratio, and then center cropped to $224\times 224$. For CIFAR-10 experiments, we initialize the classifier with ResNet-50 pre-trained on ImageNet and also provide the results trained from scratch (see Appendix C.2). For ImageNet, we train ResNet-50 from scratch.
Diffusion model. For CIFAR-10 generation, we consider two diffusion models: EDM (Karras et al., 2022) and U-ViT (Bao et al., 2023). For ImageNet-256, we explore three diffusion models: EDM2 (Karras et al., 2023), U-ViT-H/2 (Bao et al., 2023), and DiT-XL/2 (Peebles & Xie, 2023). Since EDM2 is originally designed for ImageNet-512 generation, we resize its generated images from 512 to 256 resolution, with EDM2-XXL achieving an FID of 2.14, comparable to U-ViT-H/2 and DiT-XL/2. For ImageNet-512 generation, we use EDM2-XXL (Karras et al., 2023). For COCO generation, we consider three text-to-image diffusion models: Pixart-$\alpha$ (Chen et al., 2023), SDXL (Podell et al., 2023), and Playground-v2.5 (Li et al., 2024).
@@
5 Generated distributions are easily classified as generated
In this section, we explore how the distribution generated by diffusion models, known for their success in image generation, differs from the real distribution. To analyze this, we use a distribution classification task to assess the extent of the difference.
Table 1: Binary distribution classification on the same distribution. Classifiers are indeed unable to distinguish samples from the same distribution, initially demonstrating the validity of our approach.
Distribution1 Distribution2 Classifier Accuracy (%)
U (Bao et al., 2023) U ResNet-50 50.83
ViT-S 50.57
ConvNeXt-T 50.18
E (Karras et al., 2022) E ResNet-50 50.67
ViT-S 50.44
ConvNeXt-T 50.14
5.1 Classifiers cannot distinguish samples from the same distribution
First, it is crucial to evaluate whether classifiers are indeed unable to distinguish samples from the same distribution. As shown in Tab. 1, we report the classification accuracy for the same distributions in the label-to-image scenario. Our findings indicate that classifiers are unable to distinguish samples from the same distribution, initially demonstrating the validity of the distribution classification task.
@@ 5.2 Classifiers achieve high accuracy across various settings
Label-to-image. As shown in Tab. 2, we report the binary distribution classification accuracy for various combinations of real and generated distributions in the label-to-image scenario. Surprisingly, despite the low FID scores achieved by diffusion models, neural network-based classifiers still identify significant differences between the generated and real distributions. Specifically, all classifiers easily distinguish real from generated images across different dataset combinations, with most achieving over 98% accuracy in distribution classification. Moreover, we observe a positive correlation between distribution classification accuracy and the number of training samples for the combination of U-ViT-H/2 and ImageNet-256 (see Fig. 2). This suggests that as sample size increases, classifiers learn more generalizable features rather than merely memorizing the data. Notably, with only 50k samples per distribution (less than 10% of the ImageNet samples, approximately 1.28 million, used to train the generator), classifiers achieve over 99.5% accuracy. Besides, we extend our evaluation to other generative models, e.g., GANs (see detailed results in Sec. C.1).
Table 2: Binary distribution classification on label-to-image. All classifiers yield high accuracy on various combinations of datasets and generative models. FIDs are taken from related references.
@@
Table 3: Four-way distribution classification on text-to-image. All classifiers yield high accuracy in distinguishing four distributions: COCO (Lin et al., 2014), Pixart-$\alpha$ (Chen et al., 2023), SDXL (Podell et al., 2023), and Playground-v2.5 (Li et al., 2024), using only 5k or 10k training samples per dataset. Notably, as the number of training samples increases, the accuracy of distribution classification consistently improves.
@@
Figure 2: Binary distribution classification on label-to-image. A positive correlation is observed between accuracy and the number of training samples. With only 50k samples per distribution, classifiers achieve over 99.5% accuracy.
Text-to-image. As shown in Tab. 3, we report the distribution classification accuracy in the text-to-image scenario. We consider combinations of four distributions: COCO (Lin et al., 2014), Pixart-$\alpha$ (Chen et al., 2023), SDXL (Podell et al., 2023), and Playground-v2.5 (Li et al., 2024). In this task, classifiers are not only required to distinguish between real and generated distributions but also correctly identify the specific generated distribution an image comes from. Remarkably, classifiers achieve an accuracy of over 76% across all settings. Furthermore, we observe that increasing the training sample size from 5k to 10k results in significant accuracy gains, consistent with the patterns seen in the label-to-image scenario.
5.3 Self-supervised classifiers can also identify generated images
In addition to the supervised learning protocol, we also explore self-supervised learning. Specifically, we use self-supervised pre-trained models, such as MAE (He et al., 2022) and MoCo v3 (Chen et al., 2021), to extract image features followed by training a linear classifier on these features, referred to as the self-supervised classifier. The distribution classification accuracy of these classifiers is in Tab. 4. Notably, the self-supervised classifiers also demonstrate a strong ability to distinguish between real and generated distributions. Moreover, we observed two intriguing phenomena: First, unlike in semantic classification tasks on ImageNet, the masked image modeling-based MAE outperforms the contrastive learning-based MoCo v3 in identifying generated images related to ImageNet. Second, while self-supervised classifiers achieve relatively high accuracy on ImageNet-related tasks, their performance drops significantly when applied to CIFAR-10. We hypothesize this decline can be contributed to two factors: (1) MAE and MoCo v3 are pre-trained on ImageNet, not CIFAR-10, and (2) diffusion models perform exceptionally well on CIFAR-10 image generation, making it challenging for a simple linear classifier to differentiate between real and generated images. Besides, we also evaluate self-supervised classifiers in text-to-image scenarios (see detailed results in Sec. C.3).
Table 4: Self-supervised classifiers demonstrate the ability to distinguish between real and generated distributions. For each self-supervised method, we use ViT-B as the backbone.